# Modality Translation for Object Detection Adaptation Without Forgetting Prior Knowledge

### v1

@@ Abstract
A common practice in deep learning consists of training large neural networks on massive datasets to perform accurately for different domains and tasks. While this methodology may work well in numerous application areas, it only applies across modalities due to a larger distribution shift in data captured using different sensors. This paper focuses on the problem of adapting a large object detection model to one or multiple modalities while being efficient. To do so, we propose ModTr as an alternative to the common approach of fine-tuning large models. ModTr consists of adapting the input with a small transformation network trained to minimize the detection loss directly. The original model can therefore work on the translated inputs without any further change or fine-tuning to its parameters. Experimental results on translating from IR to RGB images on two well-known datasets show that this simple ModTr approach provides detectors that can perform comparably or better than the standard fine-tuning without forgetting the original knowledge. This opens the doors to a more flexible and efficient service-based detection pipeline in which, instead of using a different detector for each modality, a unique and unaltered server is constantly running, where multiple modalities with the corresponding translations can query it. Code: https://github.com/heitorrapela/ModTr.
@@
Figure 1: Bounding box predictions over different OD methods for infrared images on two benchmarks: LLVIP and FLIR. Yellow and red boxes show the ground truth and predicted detections, respectively. FastCUT is an unsupervised image translation approach that takes as input infrared images (IR) and produces pseudo-RGB images. It does not focus on detection and requires both modalities for training. Fine-tuning is the standard approach to adapting the detector to the new modality. It requires only IR data but forgets the original knowledge of the pre-trained detector. Finally, ModTr, our approach focuses the translation on detection, requires only IR data, and does not forget the original knowledge so that it can be reused for other tasks.
@@ 1 Introduction
These additional modalities, though not as common as RGB images, are still important in various tasks, like surveillance [5], autonomous driving [35], and robotics [33], which strive to achieve robust performance across environmental changes such as different illumination conditions [2]. The dominant way to adapt pre-trained detectors to these novel modalities is by fine-tuning the model. However, fine-tuning often results in catastrophic forgetting and can destroy the intrinsic knowledge of the detector [20]. Ideally, we would like to adapt the detector to new modalities without changing the original model. This is most useful for server-side applications, where a single model runs uninterrupted and different inputs, ideally on different modalities, can query it. The main challenge here is the significant distribution shift between the new modalities, such as the difference between the visual information in RGB images and the thermal data in IR images. This shift can degrade the performance of models when applied directly to new input, as the features learned from one modality may not be relevant or present in another. This can ultimately impact the final performance of the detector [38].
@@
Image translation methods have emerged as a powerful tool to overcome the downsides of fine-tuning and narrowing the gap between source and target modalities [14]. These methods do not directly work on the weight space of the original detector but rather adapt the input values to reduce the discrepancy between the source and target modalities. However, such methods often require access to source data or some statistics about it during training. Furthermore, their primary focus is on image reconstruction quality rather than the final detection task, which can cause a significant drop in performance. For instance, in Figure 1, some detections (in red) using these generated images, in b) and d), and the fine-tuning in c) are illustrated. Additionally, results for different methods are provided in supplementary material.
(a) N-Detectors
(b) N-ModTr-1-Detector
(c) 1-Detector
@@
Figure 2: Different approaches to deal with multiple modalities and/or domains. (a) The simplest approach is to use a different detector adapted to each modality. This can lead to a high level of accuracy but requires storing models in memory multiple. (b) Our proposed solution is based on using a single pre-trained model normally trained on the more abundant data (RGB) and then adapting the input through our ModTr model. (c) A single detector is trained on all modalities jointly. This allows using of a single model but requires access to all modalities jointly, which is often not possible, especially when dealing with large pre-trained models.
@@
Our work aims to improve the image translation paradigm while addressing its limitations. Our proposed approach, Modality Translation for OD (ModTr), incorporates the detector’s knowledge into the translation module by training directly for the final detection task. Unlike traditional image translation methods, ModTr doesn’t require any source data. It is a simple approach that can be easily integrated with any detector, be it a one-stage or two-stage detector. This approach also has applications in many settings. By incorporating new modalities for a pre-trained detector, it can act as a server that receives diverse images translated from each ModTr block trained for each modality. The detector then generates the desired output with comparable performance to fully fine-tuning but requiring much less computational power on the entire system. In Figure 2, we present several options to integrate new modalities into our system. Figure 2(a) illustrates the N-Detectors approach, where each detector is trained on a specific modality. While this method is effective, it requires a significant amount of computational resources. Alternatively, we have the N-ModTr-1-Detector approach in Figure 2(b), which is less resource-intensive as it involves training a specialized model for each modality. Finally, Figure 2(c) shows a single detector trained on joint modalities, which is not memory-intensive. However, it may not be as optimal as the other methods due to the challenges in multimodal learning. In this work, we focus on the effectiveness of our approach for IR modality, which is commonly used in surveillance and robotics, and the incremental modality detector server-based application, which are important for many settings that require non-interrupt detection predictions.
Our main contributions can be summarized as follows:
@@
(3) Our proposed approach, ModTr, is evaluated in several scenarios, showcasing its advantages and flexibility. In particular, ModTr achieves competitive OD accuracy compared with image translation methods on two challenging visible/infrared datasets (LLVIP and FLIR).
@@ 2 Related Work
Object detection. OD is a computer vision task that has as its objective to provide labels and localization for the objects in the image [41]. The two categories of OD are two-stage and one-stage detectors. Two-stage detectors, exemplified by Faster R-CNN [34], first generate regions of interest and then use a second classifier to confirm object presence within those regions. On the other hand, one-stage detectors streamline the detection process by eliminating the proposal generation stage, aiming for end-to-end training and real-time inference speeds. RetinaNet [23] is a one-stage OD model that utilizes a focal loss function to address class imbalance during training. Also, models like FCOS [36] have emerged in this category, eliminating predefined anchor boxes to potentially enhance inference efficiency. The proposed work investigates these three traditional and powerful detectors: the two-stage Faster R-CNN detector, which is highly used on the benchmarks, RetinaNet, and FCOS. The choice of such detectors was due to the simplicity in implementation and integration among other methods, as well as a different range of pre-trained backbone weights, such as ResNet and MobileNet.
@@
Image Translation. Image translation is a pivotal task in computer vision, aiming to map images from a source domain to a target domain while preserving inherent content [30]. The goal is to discover a transformation function such that the distribution of images in the translated domain is aligned with the distribution of images in the target domain. The commonly used approaches for image translation are based on variational autoencoders (VAEs) [19] and generative adversarial network (GANs) [10, 30]. Isola et al. developed the Pix2Pix [17], a method that consists of a generator (based on U-Net) and a discriminator (based on GANs architecture) that work together to generate images based on input data and labels. Then, Zhu et al. proposed a method called CycleGAN [44], which is based on GANs, with the objective of unsupervised domain translation. Even though CycleGAN can produce quite visual results, it’s hard to optimize due to the adversarial mechanism and memory footprint needed. In contrast, VAEs are easier to train than GANs but require more constraints in the optimization to produce images of good quality than GAN-based approaches. Recent advancements include diffusion models known for their high-quality image generation, although they may not inherently suit domain translation tasks. To enhance models such as CycleGAN, novel methods like Contrastive Unpaired Translation (CUT) [31] and FastCUT [31] have been introduced. CUT, in particular, accelerates the image translation process by maximizing mutual information between image patches, achieving competitive results quickly. In the context of RGB/IR modality, InfraGAN presents an image-level adaptation for RGB to IR conversion, prioritizing image quality [29]. This approach is distinct in its focus on optimizing image quality losses. Moreover, Herrmann et al. have explored OD in RGB/IR modality by adapting IR images to RGB using traditional image preprocessing techniques, allowing the use of RGB object detectors without parameter modification [12]. Recently, there have been many advances in image translation, but they do not target OD tasks. Therefore, Medeiros et al. proposed the HalluciDet [25], which uses an image translation mechanism for OD, but in their work, they have as a prior the access to the source data to pre-train the detector on the same RGB domain of the OD modality that the model needs to adapt.
@@
Adapting without forgetting. Catastrophic forgetting (CF) is the idea that a neural network tends to forget knowledge when sequentially trained on a different task and replaces it with knowledge tailored to the new objective [39]. CF can be harmful or beneficial. Researchers identified harmful learning as situations where retaining the original knowledge while adapting to a different task is necessary. In that case, it is imperative to mitigate the risk of CF. However, some CF can also be beneficial, for instance, to prevent privacy leakage from large pre-trained models, to enhance the generalization, or to remove noisy information from the originally, acquired knowledge that is negatively affecting the new tasks. In our case, knowledge-forgetting is harmful. There are different ways to address this issue including simple techniques like decreasing the learning rate [13], use weight decay [4, 43] or mixout regularization [21] during fine-tuning or more complex approaches like Recall and learn [6], Robust Information Fine-tuning [40] or CoSDA [9]. Some adaptation methods use techniques based on replay of the source data or even using the weights of the initial model to keep some prior information [26]. However, these methods may still produce a loss of knowledge since the original parameters are not frozen. Furthermore, in adapting without forgetting, an Adapter, which adopts a frozen pre-trained backbone to generate a representation followed by a different classifier for each downstream task [39], can be seen as a powerful method for keeping knowledge. Even though our ModTr shares some similarities, we work in the input space to adapt to the new modalities, and we address this incremental modality adaptation, optimizing the translation directly for the final OD task.
@@ 3 Proposed Method
Preliminary Definitions. We denote the training set for the OD task as $\mathcal{D}=\{(x,\mathcal{Y})\}$, where $x\in\mathbb{R}^{W\times H\times C}$ represents an image in the dataset, with dimensions $W\times H$ and $C$ channels. Subsequently, the object detection model aims to identify $N$ regions of interest within these images, denoted as $\mathcal{Y}=\{(b_{i},c_{i})\}_{i=1}^{N}$. Each region of interest $b_{i}$ is defined by the top-left corner coordinates and the width and height of the object. Additionally, a classification label $c_{i}$ is assigned to each detected object, indicating its corresponding class within the dataset. In this study, the number of input channels for the detector is fixed at three, corresponding to RGB-like inputs. In terms of optimization, the primary goal of this task is to maximize detection accuracy, often measured using the average precision (AP) metric across all classes. An OD is formally represented as the mapping $f_{\theta}:\mathbb{R}^{W\times H\times C}\to\mathcal{Y}$, where $\theta$ denotes the parameter vector. To train such a detector effectively, a differentiable surrogate for the AP metric, referred to as the detection cost function, $\mathcal{C}_{det}(\theta)$, is employed. The typical structure of such a cost function involves computing the average detection loss over dataset $\mathcal{D}$, denoted as $\mathcal{L}_{det}$:
$\begin{split}\mathcal{L}_{det}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{(x,\textbf{y})\in\mathcal{D}}\mathcal{L}_{det}[f_{\theta}(x),\mathcal{Y}].\end{split}$ (1)
@@
Modality Translation Module (ModTr). Our approach primarily consists of an image-to-image translation network responsible for converting the input modality into an RGB-like space intelligible to the detector. These networks typically adopt an encoder-decoder structure to synthesize and reconstruct knowledge in a pixel-wise manner. While we employ U-Net as the translation network in this work, our framework is general and not limited by the translation architecture. In general terms, this mapping is denoted as $h_{\vartheta}^{d}\colon\mathbb{R}^{W\times H\times C}\to\mathbb{R}^{W\times H\times 3}$, with a translation network assigned to each available input modality $d$. Unlike the detection network, the number of input channels varies depending on the modality, for instance, $C=1$ for IR and depth images. It’s important to note that, being a pixel-level architecture, the output of such a network retains the spatial resolution of the input. However, the number of output channels is consistently fixed at three, corresponding to RGB-like images.
@@
Unlike other image-to-image translation approaches, we drive the translation process using the aforementioned detection cost. Thus, the underlying optimization problem is formulated as $\vartheta^{*}=\arg\min\mathcal{L}_{det}(\vartheta)$, incorporating the output of the composition $(f_{\theta}\circ h_{\vartheta}^{d})(x)$ at the loss function level. To streamline the learning process, we utilize a residual learning strategy in which the function $h_{\vartheta}^{i}$ focuses on capturing the small variations in the input that are necessary to solve the task. This approach is similar to the one employed on diffusion models, which served as inspiration for our work. For the sake of simplicity, we separate the fusion step from the translation mapping in our notation, as various types of fusion are investigated. Consequently, the proposed image-to-image translation loss function is defined as:
@@
$\begin{split}\mathcal{L}_{\text{ModTr}}(x,\mathcal{Y};\vartheta)&=\mathcal{L}_{det}[f_{\theta}\left(\Phi(h_{\vartheta}^{d}(x),x)\right),\mathcal{Y}],\end{split}$ (2)
@@
In addition, note that, while a detection loss is employed to update the translation network, the weight vector $\theta$ remains constant. This constraint is consistent with the premise of this study, where a pre-trained detector is solely available on the server side and remains unaltered. An overview of the proposed approach can be seen in Fig. 2(b).
Fusion strategy. As previously mentioned, we utilize a non-parametric fusion of the intermediate representation $h_{\vartheta}(x)$ and the original input $x$ to simplify the learning process of the translation network. In this context, we investigate three straightforward non-parametric fusion functions inspired by previous works. The first approach involves element-wise summation, commonly used in ResNet and diffusion models. The second approach employs an element-wise product, also known as the Hadamard product, which is particularly interesting for attention mechanisms and has been explored previously for re-calibrating feature maps based on their importance [15]. The final approach is based on DenseFuse [22]; however, instead of applying it at the feature level, we apply it between the input and output of the translation network.
a) ModTr+: The addition mechanism involves forwarding the input modality and summing it with the output of the translation network. This residual connection of the input serves as regularization for the model, aiding the network in learning the missing information necessary for detector operation. This operator learns the new representation by amplifying pixel values when the weights of the translation representation tend toward $1$, or preserving the original information when they tend toward 00. Such a range of values is due to our modification on the U-Net to generate, in which we changed the last layer to a sigmoid layer so we can better control the generated image to be closer to a real RGB-like image:
$\begin{split}\mathcal{L}_{\text{ModTr}_{+}}(x,\mathcal{Y};\vartheta)&=\mathcal{L}_{det}[f_{\theta}\left(h_{\vartheta}^{d}(x)+x\right),\mathcal{Y}].\end{split}$ (3)
b) ModTr: The Hadamard product-based fusion serves as a gating mechanism to filter or highlight information from the input image. In this approach, the output of the translation network acts as a weight map for the input, and they are fused using pixel-wise multiplication, $\odot$. Consequently, the translation networks tend to highlight information from the input when the pixel value tends toward $1$ or discard it when it approaches 00. Additionally, the output translation modality can be interpreted as an attention map:
$\begin{split}\mathcal{L}_{\text{ModTr}_{\odot}}(x,\mathcal{Y};\vartheta)&=\mathcal{L}_{det}[f_{\theta}\left(h_{\vartheta}^{d}(x)\odot x\right),\mathcal{Y}].\end{split}$ (4)
c) ModTr: The subsequent fusion mechanism draws inspiration from DenseFuse [22], which employs the relative importance of pixels as an attention mechanism for both the translation network’s output and input. This attention mechanism operates by providing a weighted average of the channels. The implementation details for such an operator can be found on [22]. Then, the proposed loss function is given by:
$\begin{split}\mathcal{L}_{\text{ModTr}_{\oplus}}(x,\mathcal{Y};\vartheta)&=\mathcal{L}_{det}[f_{\theta}\left(\Phi(h_{\vartheta}^{d}(x),x)\right),\mathcal{Y}],\end{split}$ (5)
with
$\Phi(x,\hat{x})=\frac{x\odot e^{x}+\hat{x}\odot e^{\hat{x}}}{e^{x}+e^{\hat{x}}}.$

### v3

@@ Abstract
A common practice in deep learning involves training large neural networks on massive datasets to achieve high accuracy across various domains and tasks. While this approach works well in many application areas, it often fails drastically when processing data from a new modality with a significant distribution shift from the data used to pre-train the model. This paper focuses on adapting a large object detection model trained on RGB images to new data extracted from IR images with a substantial modality shift. We propose Modality Translator (ModTr) as an alternative to the common approach of fine-tuning a large model to the new modality. ModTr adapts the IR input image with a small transformation network trained to directly minimize the detection loss. The original RGB model can then work on the translated inputs without any further changes or fine-tuning to its parameters. Experimental results on translating from IR to RGB images on two well-known datasets show that our simple approach provides detectors that perform comparably or better than standard fine-tuning, without forgetting the knowledge of the original model. This opens the door to a more flexible and efficient service-based detection pipeline, where a unique and unaltered server, such as an RGB detector, runs constantly while being queried by different modalities, such as IR with the corresponding translations model. Our code is available at: https://github.com/heitorrapela/ModTr.
@@
Figure 1: Bounding box predictions over different adaptations of the RGB detector (Faster R-CNN) for IR images on two benchmarks: LLVIP and FLIR. Yellow and red boxes show the ground truth and predicted detections, respectively. In a) we see the RGB data. In b) FastCUT is an unsupervised image translation approach that takes as input infrared images (IR) and produces pseudo-RGB images. It does not focus on detection and requires both modalities for training. In c) we have fine-tuning, which is the standard approach to adapting the detector to the new modality. It requires only IR data but forgets the original knowledge of the original RGB detector. Finally, in d) is the ModTr, which focuses the translation on detection, requires only IR data and does not forget the original knowledge so that it can be reused for other tasks. Bounding box predictions for other detectors are provided in the supplementary material.
@@ 1 Introduction
These additional modalities, though not as common as RGB images, are still important in various tasks, like surveillance [6, 9], autonomous driving [43, 35], and robotics [40, 24], which strive to achieve robust performance in real-world environments, where capture conditions change, such as different illumination conditions [3]. The dominant way to adapt pre-trained detectors to these novel conditions is by fine-tuning the model [31]. However, fine-tuning often results in catastrophic forgetting and can destroy the intrinsic knowledge of the detector [26]. Ideally, we would like to adapt the detector to new modalities without changing the original model. This is most useful for server-side applications, where a single model runs uninterrupted and can be queried by different inputs, ideally on different modalities. The main challenge lies in the significant distribution shift introduced by the new modality. This shift occurs because the pre-trained knowledge, such as the visual information in RGB images, differs markedly from the thermal data in IR images. This shift can degrade model performance when applied directly as input to the model, since the features learned from one modality may not be relevant or present in another. This can ultimately impact the resulting OD performance [46].
@@
Image translation methods [37, 38] have emerged as powerful tools to overcome the downsides of fine-tuning and narrowing the gap between source and target modalities [18]. These methods do not directly work on the weight space of the original detector but rather adapt the input values to reduce the discrepancy between the source and target modalities. However, such methods often require access to source data or some statistics about it during training. Furthermore, their primary focus is on image reconstruction quality rather than the final OD task, which can cause a significant drop in performance. For instance,  Figure 1 shows different ways to adapt the RGB detector (see the caption for more details).
@@
Figure 2: Different approaches to deal with multiple modalities and/or domains. (a) The simplest approach is to use a different detector adapted to each modality. This can lead to a high level of accuracy but requires storing several models in memory. (b) Our proposed solution uses a single pre-trained model normally trained on the more abundant data (RGB) and then adapts the input through our ModTr model. (c) A single detector is jointly trained on all modalities. This allows using of a single model but requires access to all modalities jointly, which is often impossible, especially when dealing with large pre-trained models.
@@
Our work aims to improve the image translation paradigm while addressing its limitations. Our proposed approach, Modality Translation for OD (ModTr), incorporates the detector’s knowledge into the translation module by training directly for the final detection task. Unlike traditional image translation methods, ModTr does not require any source data. It is a conceptually simple approach that can be easily integrated with any detector, be it a one-stage or two-stage detector. A notable application of ModTr is using a pre-trained RGB detector as a server that incorporates different ModTr blocks as input translators for new modalities such as IR. This new detector generates the desired output with performance comparable to full fine-tuning without losing the original knowledge of the pre-trained model. In Fig. 2, we present several options for integrating IR modalities into an RGB system. Fig. 2a illustrates the N-Detectors approach, where each detector is trained for a specific case. This method effectively demands more memory and forgets previously learned information. Fig. 2c shows a single detector trained on combined modalities. This method does not incur additional memory, yet it requires simultaneous access to all modalities, which may not always be feasible. Fig. 2b illustrates our proposed approach, which involves training a specialized translator for each condition without altering the parameters of the original detector. The N-ModTr-1-Detector strikes a balance between the previous methods, addressing their shortcomings by requiring only a single detector. Importantly, it retains the original pre-training knowledge, as it leaves the detector unchanged. In this work, we focus on the effectiveness of our approach for the IR modality, commonly used in surveillance and robotics, and the incremental modality detector server-based application, which is crucial for many settings that require uninterrupted detection predictions.
Our main contributions can be summarized as follows.
@@
(3) An extensive empirical evaluation of ModTr in several scenarios, showcasing its advantages and flexibility. In particular, with our different proposed fusion strategies, ModTr achieves OD accuracy that is competitive when compared with image translation methods on two challenging RGB/IR datasets (LLVIP and FLIR).
@@ 2 Related Work
(a) Object Detection. OD is a computer vision task that aims to provide labels and localization for the objects in the image [49]. Two-stage detectors, exemplified by Faster R-CNN [41], generate regions of interest and then use a second classifier to confirm object presence within those regions. On the other hand, one-stage detectors streamline the detection process by eliminating the proposal generation stage, aiming for end-to-end training and real-time inference speeds. RetinaNet [29] is a one-stage OD model that utilizes a focal loss function to address class imbalance during training. Also, models like FCOS [44] have emerged in this category, eliminating predefined anchor boxes to potentially enhance inference efficiency. The proposed work investigates these three traditional and powerful detectors: Faster R-CNN, RetinaNet, and FCOS. The choice of such detectors was due to the simplicity in implementation and integration among other methods, as well as a different range of pre-trained backbone weights, such as ResNet [14] and MobileNet [16].
@@
(b) Image Translation. Image translation is a pivotal task in computer vision, aiming to map images from a source domain to a target domain while preserving inherent content [37]. The goal is to discover a transformation function such that the distribution of images in the translated domain is aligned with the distribution of images in the target domain. The commonly used approaches for image translation are based on variational autoencoders (VAEs) [25] and generative adversarial network (GANs) [12, 37]. Isola et al. developed the Pix2Pix [22], a method that consists of a generator (based on U-Net) and a discriminator (based on GANs architecture) that work together to generate images based on input data and labels. Then, Zhu et al. proposed a method called CycleGAN [52], which is based on GANs, with the objective of unsupervised domain translation. Even though CycleGAN can produce quite visual results, it’s hard to optimize due to the adversarial mechanism and memory footprint needed. In contrast, VAEs are easier to train than GANs but require more constraints in the optimization to produce images of good quality than GAN-based approaches. Recent advancements include diffusion models known for their high-quality image generation, although they may not inherently suit domain translation tasks. To enhance models such as CycleGAN, novel methods like Contrastive Unpaired Translation (CUT) [38] and FastCUT [38] have been introduced. CUT, in particular, accelerates the image translation process by maximizing mutual information between image patches, achieving competitive results quickly. In the context of RGB/IR modality, InfraGAN presents an image-level adaptation for RGB to IR conversion, prioritizing image quality [36]. This approach is distinct in its focus on optimizing image quality losses. Moreover, Herrmann et al. have explored OD in RGB/IR modality by adapting IR images to RGB using traditional image preprocessing techniques, allowing the use of RGB object detectors without parameter modification [15]. Despite significant advances in image translation, these techniques do not specifically address OD tasks. In our previous work, we introduced HalluciDet [31], which employs an image translation mechanism for OD. However, this approach requires prior access to the source RGB data from the same domain as the target for pre-training the detector.
@@
(c) Adapting Without Forgetting. Catastrophic forgetting (CF) is the idea that a neural network tends to forget knowledge when sequentially trained on a different task and replaces it with knowledge tailored to the new objective [47]. CF can be harmful or beneficial. Researchers identified harmful learning as situations where retaining the original knowledge while adapting to a different task is necessary. In that case, it is imperative to mitigate the risk of CF. However, some CF can also be beneficial, for instance, to prevent privacy leakage from large pre-trained models, to enhance the generalization, or to remove noisy information from the originally, acquired knowledge that is negatively affecting the new tasks. In our case, knowledge-forgetting is harmful. There are different ways to address this issue including simple techniques like decreasing the learning rate [17], use weight decay [5, 51] or mixout regularization [27] during fine-tuning or more complex approaches like Recall and learn [7], Robust Information Fine-tuning [48] or CoSDA [11]. Some adaptation methods use techniques based on replay of the source data or even using the weights of the initial model to keep some prior information [32]. Some of these works focus on adding continually different tasks in an incremental learning setting. However, these methods may still produce a loss of knowledge since the original parameters are not frozen. Furthermore, in adapting without forgetting, an adapter, which adopts a frozen pre-trained backbone to generate a representation followed by a different classifier for each downstream task [47], can be seen as a powerful method to preserve knowledge. Even though our ModTr shares some similarities, we work in the input space to adapt to the new modalities, and address this incremental modality adaptation, optimizing the translation directly for the final OD task.
@@ 3 Proposed Method
(a) Preliminary Definitions. The training set for OD is denoted as $\mathcal{D}=\{(x,Y)\}$, where $x\in\mathbb{R}^{W\times H\times C}$ represents an image in the dataset, with dimensions $W\times H$ and $C$ channels. Subsequently, the OD model aims to identify $N$ regions of interest within these images, denoted as $Y=\{(b_{i},c_{i})\}_{i=1}^{N}$. The top-left corner coordinates and the width and height of the object define each region of interest $b_{i}$. Additionally, a classification label $c_{i}$ is assigned to each detected object, indicating its corresponding class within the dataset. In this study, the number of input channels for the detector is fixed at three, corresponding to RGB-like inputs. In terms of optimization, the primary goal of this task is to maximize detection accuracy, often measured using the average precision (AP) metric across all classes. An OD is formally represented as the mapping $f_{\theta}:\mathbb{R}^{W\times H\times C}\to\hat{Y}$, where $\theta$ denotes the parameter vector. To effectively train a detector, a differentiable surrogate for the AP metric, referred to as the detection cost function, $\mathcal{C}_{det}(\theta)$, is employed. The typical structure of such a cost function involves computing the average detection loss over dataset $\mathcal{D}$, denoted as $\mathcal{L}_{det}$, described as:
$\begin{split}\mathcal{C}_{det}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{(x,Y)\in\mathcal{D}}\mathcal{L}_{det}(f_{\theta}(x),Y).\end{split}$ (1)
@@
(b) Modality Translation Module. Our approach primarily consists of an image-to-image translation network responsible for converting the input modality into an RGB-like space intelligible to the detector. These networks typically adopt an encoder-decoder structure to synthesize and reconstruct knowledge in a pixel-wise manner. While we employ U-Net [42] as the translation network, with parameters $\vartheta$, in this work, our framework is general and not limited by the translation architecture. In general terms, this mapping is denoted as $h_{\vartheta}^{d}\colon\mathbb{R}^{W\times H\times C}\to\mathbb{R}^{W\times H\times 3}$, with a translation network assigned to each available input modality $d$. Unlike the detection network, the number of input channels varies depending on the modality, for instance, $C=1$ for IR and depth images. It’s important to note that, being a pixel-level architecture, the output of such a network retains the spatial resolution of the input. However, the number of output channels is consistently fixed at three, corresponding to RGB-like images ($C=3$).
@@
Unlike other image-to-image translation approaches, we drive the process using the aforementioned detection cost (Equation (1)). Thus, the underlying optimization problem is formulated as $\vartheta^{*}=\arg\min\mathcal{L}_{det}(\vartheta)$, incorporating the output of the composition $(f_{\theta}\circ h_{\vartheta}^{d})(x)$ at the loss function level. To streamline the learning process, we utilize a residual learning strategy in which the function $h_{\vartheta}^{d}$ focuses on capturing the small variations in the input that are necessary to solve the task. This approach is similar to the one employed on diffusion models, which inspired our work. For the sake of simplicity, we separate the fusion step from the translation mapping in our notation, as various types of fusion are investigated. Consequently, the proposed image-to-image translation loss function is defined as:
@@
$\begin{split}\mathcal{L}_{\text{ModTr}}(x,Y;\vartheta)&=\mathcal{L}_{det}(f_{\theta}\left(\Phi(h_{\vartheta}^{d}(x),x)\right),Y),\end{split}$ (2)
@@
In addition, note that, while a detection loss is employed to update the translation network, the weight vector $\theta$ remains constant. This constraint is consistent with the premise of this study, where a pre-trained detector is solely available on the server side and remains unaltered. An overview of the proposed approach can be seen in Fig. 2 b).
(c) Fusion strategy. As previously mentioned, we utilize a non-parametric fusion of the intermediate representation $h_{\vartheta}^{d}(x)$ and the original input $x$ to simplify the learning process of the translation network. In this context, we employ an element-wise product, also known as the Hadamard product, which is particularly interesting for attention mechanisms and has been explored previously for re-calibrating feature maps based on their importance [20]. Although we investigated various fusion mechanisms, the element-wise product yielded the best results. For more details on different fusion strategies, please refer to supplementary materials.
ModTr: The Hadamard product-based fusion serves as a gating mechanism to filter or highlight information from the input image. In this approach, the output of the translation network acts as a weight map for the input, and they are fused using pixel-wise multiplication, $\odot$. Consequently, the translation network tends to highlight information from the input when the pixel value tends toward $1$ or discard it when it approaches 00. Additionally, the output translation modality can be interpreted as an attention map, as described by the following Equation (3):
$\begin{split}\mathcal{L}_{\text{ModTr}_{\odot}}(x,Y;\vartheta)&=\mathcal{L}_{det}(f_{\theta}\left(h_{\vartheta}^{d}(x)\odot x\right),Y).\end{split}$ (3)